OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 8 and 15 units in length.

Parth (parthkohli):

Try using the Pythagorean theorem.

OpenStudy (callisto):

Use Pyth Thm \[hypotenuse = \sqrt{8^2+15^2} = ...?\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow a^2 + b^2 = c^2}\) Since c is the hypotenuse, then substitute a and b with 8 and 15 in any order.

OpenStudy (anonymous):

Pythagoras theorem says : \[(Hypotenuse)^2 = (Base)^2 + (Height)^2\] So,\[H^2 = 8^2 + 15^2 = 64 + 225\] \[H^2 = 289\] Taking square root both the sides, \[H = \sqrt{289}\] Hypotenuse Length = 17 Units is the answser...

OpenStudy (saranya):

a^2+b^2=c^2 8^2+15^2=289 sq root of 289=17

OpenStudy (anonymous):

thanks guys!